At a temperature of 20 degree centigrade the common amoeba reproduces by splitting in half every 24 hours. if we start with a single amoeba how many will there be after (a) 8 days (b) 16 days?
Answer by venugopalramana(3286)

At a temperature of 20 degree centigrade the common amoeba reproduces by splitting in half every 24 hours.
THAT IS IN 1 DAY 1 BECOMES 2..IN 2 DAYS 2 BECOME 4...ETC...
THIS IS A G.P....1,2,4,8,16.....
ETC WITH R=2 AND A=1
SO TN=A*(R)^(N-1)....N=8...
SO T8=1*2^*(8-1)=2^7=128...SO AFTER 8 DAYS WE GET 128*2=256
T16=1*2^(16-1)=2^15=32768...SO AFTER 16 DAYS WE GET 32768*2=65536
if we start with a single amoeba how many will there be after (a) 8 days (b) 16 days?